# Robot-R1: Reinforcement Learning for Enhanced Embodied Reasoning in Robotics

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have recently shown great promise in advancing robotics by combining embodied reasoning with robot control. A common approach involves training on embodied reasoning tasks related to robot control using Supervised Fine-Tuning (SFT). However, SFT datasets are often heuristically constructed and not explicitly optimized for improving robot control. Furthermore, SFT often leads to issues such as catastrophic forgetting and reduced generalization performance. To address these limitations, we introduce Robot-R1, a novel framework that leverages reinforcement learning to enhance embodied reasoning specifically for robot control. Robot-R1 learns to predict the next keypoint state required for task completion, conditioned on the current scene image and environment metadata derived from expert demonstrations. Inspired by the DeepSeek-R1 learning approach, Robot-R1 samples reasoning-based responses and reinforces those that lead to more accurate predictions. To rigorously evaluate Robot-R1, we also introduce a new benchmark that demands the diverse embodied reasoning capabilities for the task. Our experiments show that models trained with Robot-R1 outperform SFT methods on embodied reasoning tasks. Despite having only 7B parameters, Robot-R1 even surpasses GPT-4o on reasoning tasks related to low-level action control, such as spatial and primitive movement reasoning.
**footnotetext: Equal advising.
@@ 1 Introduction
Figure 1: Illustration of the Robot-R1 framework. (a) Robot-R1 uses robot states and image observations from expert demonstrations to create a dataset. (b) These data are reformulated into three different multiple-choice question answering (MCQA) tasks: predicting next states, current states, and primitive movements. (c) During training, an LVLM solves MCQA tasks with reasoning which is then optimized using the GRPO algorithm [17] to reinforce reasoning pathways.
@@
The key idea of Robot-R1 is to train LVLMs to predict the next keypoint state necessary for task completion through a reasoning process, and to optimize this reasoning via RL to maximize prediction accuracy. However, since the keypoint lies in a continuous space, the action space is vast, making exploration particularly challenging. This makes it difficult for the model to efficiently learn reasoning strategies through trial and error. To address this, Robot-R1 reformulates the next-state prediction problem as a multiple-choice question-answering (QA) problem. This discrete formulation narrows the action space, making the learning process more efficient. To further enhance state understanding, we add two auxiliary QA tasks: (i) current state prediction QA, where the model predicts the robot’s current state from visual observations, and (2) primitive movement prediction QA, where the model predicts rule-based linguistic descriptions of state changes (e.g., “move up”, “move down”).
To analyze how Robot-R1 improves embodied reasoning capabilities in trained models, we design a novel benchmark called the Robot-R1 Bench (see Figure 2). Through this benchmark, we observe that models trained with Robot-R1 achieve over a 28% improvement in embodied reasoning tailored for low-level action control. In particular, despite having only the 7B parameters, Robot-R1 outperforms several major commercial models, including GPT-4o [18], in this domain. To evaluate whether the embodied reasoning abilities acquired through Robot-R1 transfer effectively to other tasks, we conduct evaluations on two external benchmarks. First, on EmbodiedBench Manipulation [19], a vision-driven robot agent benchmark, Robot-R1 yields a 31% improvement in task performance. Moreover, on SpatialRGPTbench [20], which tests 3D spatial reasoning, name achieves approximately 40% improvement in quantitative metrics and about 60% improvement in qualitative metrics. These results contrast with conventional SFT approaches, which show limited improvement in embodied reasoning and tend to exhibit performance degradation when applied to out-of-distribution tasks. This suggests that Robot-R1 can effectively learn generalizable and diverse embodied reasoning capabilities simply by next-state prediction.
@@ 3.1 Preliminaries
Group relative policy optimization. We employ the Group Relative Policy Optimization (GRPO) algorithm [17] to optimize the policy. Specifically, for a given query $q$, we generate $G$ different responses $\mathbf{o}=\{o_{1},o_{2},o_{3},,,o_{G}\}$ from the current policy $\pi_{\theta_{\text{old}}}$. Each response $o_{i}$ is evaluated using the reward model to produce a corresponding reward $r_{i}$, resulting in the set $\mathbf{r}=\{r_{1},r_{2},...r_{G}\}$. For each response, we compute an advantage score $A_{i}$, using the mean and standard deviation of the rewards for that query, i.e., $A_{i}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$. Finally, GRPO updates the policy by maximizing this advantage while applying a KL penalty, through the following objective:
@@ 3.2 Data Generation
Primitive movement prediction QA. The goal of this task is to predict the primitive movement (e.g., “move up”, “move right”) that the robot should take to move from the current state toward the next waypoint. To extract labels for these movement primitives, we apply rule-based heuristics based on the change in 3D Cartesian space. Specifically, we compute the difference in position between the current and next waypoint states, $s_{k^{*}}-s_{t}$, and identify the primary direction of movement along the $x$, $y$, and $z$ axes. If there is a non-zero change along an axis, a corresponding directional command is generated (e.g., “move forward“ for a positive change in the x-axis); this label is denoted as $a_{t}$. If the change along an axis is less than or equal to half of the largest change across all axes, the adverb “slightly” is added (e.g., “Slightly move backward”). To construct a multiple-choice question, we randomly sample three adjacent movement= $\{a^{\prime}_{d1},a^{\prime}_{d2},a^{\prime}_{d3}\}$. The question is formulated as $Q_{\tt{primitive}}(M,o_{t},s_{t},\texttt{shuffle}(\{a_{t},a^{\prime}_{d1},a^{\prime}_{d2},a^{\prime}_{d3}\}))$ where the answer is $A=a_{t}$.
@@ 3.4 Robot-R1 Bench: A Novel Benchmark for Evaluating Embodied Reasoning
Existing visual question answering (VQA) benchmarks for embodied reasoning primarily focus on general visual understanding, without explicitly evaluating the nuanced reasoning processes behind robotic behavior [48, 49, 50]. Meanwhile, robotics-focused benchmarks often rely on simulation-based evaluation [19, 51, 52], priortize task successes as the main metric, or focus on tasks that are not aligned with the table-top manipulation scenarios central to our work. While these benchmarks provide valuable insights, they do not directly assess a model’s ability to reason about robot actions. To address this gap, we introduice Robot-R1 Bench, a new benchmark designed to evaluate embodied reasoning through open-ended answering grounded in robot demonstrations.
Design features of Robot-R1 Bench. In practical applications, models typically face open-ended decisions rather than selecting from pre-defined options. To better reflect this, Robot-R1 Bench adopts an open-ended QA format that more closely resembles practical use cases. All questions are based on images from expert demonstrations, ensuring visually grounded and realistic robotic scenarios. Robot-R1 Bench supports fine-grained evaluation across four key reasoning types: high-level action planning, subtask reasoning, primitive movement reasoning, and spatial reasoning. Each question is crafted to assess one of these abilities, allowing for evaluation across both high-level decision making (action planning and subtask reasoning) and low-level control reasoning (primitive action reasoning and spatial reasoning).
Dataset construction. The Robot-R1 Bench dataset consists of 10 tasks from RLBench [53]. For each task, we randomly sample five frames from expert demonstratrions, resulting in a total of 50 images. For each image, experienced researchers manually created questions and detailed reference answers, targeting the four key reasoning abilities. The final dataset consists of 215 open-ended questions in total: 65 for spatial reasoning and 50 for each of the other three reasoning types.
@@ 4 Experiment
4.1 Experimental Setup
Table 3: Robot-R1 Bench results tailored for high-level control. Robot-R1 significantly outperforms SFT, the baseline training method.
Model Planning Subtask GPT-4o [18] 1.96 2.02 claude-3-7-sonnet [59] 1.72 1.58 gemini-2.0-flash [61] 1.84 1.12 Qwen2.5-VL-7B-Ins [62] 1.66 1.04 w/ Direct SFT 0 0.04 w/ CoT SFT 0.60 1.28 w/ Robot-R1 (Ours) 1.44 1.30
Training data. The training data used in our experiments are generated using the built-in data generator in RLBench [53]. We collect 50 demonstrations per task from the variation 0 settings, rendering them at 224 × 224 resolution using OpenGL3. Waypoints from these demonstrations are extracted using the waypoint extraction functionality provided in the ARM [47] repository [^1] https://github.com/stepjam/ARM. The tasks selected from RLBench for training our model include: pick_up_cup, push_button, put_rubbish_in_bin, phone_on_base, and take_lid_off_saucepan. For generating the waypoint prediction QA training data, as described in Section 3, a frame extraction interval of $t=10$ was used between the current frame ($o_{t},s_{t}$) and the selected future keypoint frame ($s_{k^{*}}$), similar to demonstration augmentation proposed in ARM. Consequently, each task contains approximately 2.5K questions, resulting in a total of around 7.5K QA pairs across the three QA tasks used for training. Baseline models. The proposed Robot-R1 learns to predict the next waypoint state from expert demonstrations by generating an explicit reasoning process. To evaluate the importance of this learned reasoning, we establish two baseline models trained via Supervised Fine-tuning (SFT):
@@ 4.1 Experimental Setup
    CoT SFT: This model is fine-tuned on a QA dataset where the input prompt includes a manually structured Chain-of-Thought (CoT) path. This CoT path sequentially incorporates action planning, subtask reasoning, and primitive action components to guide the prediction of the next waypoint state, similar to Zawalski et al. [15]. This baseline enables comparison between the reasoning learned via Robot-R1 against reasoning guided by a predefined, structured thought process.
@@
Implementation detail. Robot-R1 is trained using GRPO as described in Section 3. We use the hyperparameter configurations in the EasyR1 workspace. [^2] https://github.com/hiyouga/EasyR1 For the training process, we utilize a batch size of 128 over a single epoch. During GRPO updates, 5 samples were generated per prompt with a sampling temperature of 1.0. The rollout batch size is set to 512. We use a learning rate of $1.0\times 10^{-6}$ with a weight decay of $1.0\times 10^{-2}$. For the SFT baselines, we use the same batch size, but learning rate is $1.0\times 10^{-5}$. We use the hyperparameters provided in the Qwen2-VL-Finetune workspace. [^3] https://github.com/2U1/Qwen2-VL-Finetune
@@ 4.2 Robot-R1 Bench
Results. Table 2 presents the results across the reasoning benchmarks focused on low-level control, and Figure 3 provides its visualization to help understanding. We find that commercial model performance improves consistently with their version updates, indicating the benchmark’s sensitivity to model quality. Notably, the Robot-R1-trained model achieves the highest performance across both the overall benchmark and its individual components, outperforming all commercial models evaluated. Specifically, in primitive movement prediction, our method outperforms even the SFT-trained model, which is explicitly trained to predict accurate actions. Table 3 further show the performance in reasoning tasks tailored for high-level control. Here, the correlation between model version and benchmark performance is less pronounced. We hypothesize that this is due to the inherently high variability in valid high-level actions for robotic control. Nevertheless, we still observe a general trend of increasing benchmark performance as model quality improves. Importantly, our Robot-R1 exhibits significant performance gains in subtask reasoning – predicting the immediate high-level action – compared to the base and the SFT-trained model. This suggests that effective high-level reasoning can emerge from training exclusively on low-level control information, even without explicit high-level action supervision. Interestingly, we find that performance on action planning slightly decreases, likely because our training objective focuses primarily on next keypoint prediction rather than long-horizon planning.
@@ 4.3 EmbodiedBench Manipulation
+ Direct SFT 0 0 0 0 0 0
+ CoT SFT 0 0 0 0 0 0
+ Robot-R1(Ours) 12.5 8.3 6.3 14.6 16.7 11.68
@@
Setup. We evaluate the model performance on four different tasks: Pick & Place Objects, Stack Objects, Shape Sorter Placement, and Table Wiping. To comprehensively assess instruction-following capabilities, the benchmark categorizes task instructions into five groups: (i) Base, which requires fundamental task-solving skills, (ii) Common Sense, which requires general world knowledge to resolve indirect objective references, (iii) Complex instructions, which involves longer, possibly distracting context that obscures the primary command, (iv) Spatial reasoning, where objects are referred to by their spatial relationships to other objects, and (v) Visual understanding, which requires recognizing objects based on visual attributes such as color or shape. For comparison, the experimental results of GPT-4o-mini [18] and Gemini-1.5-flash [60] are adopted directly from the values reported in their paper.
@@ 4.4 Ablation Study and Analysis
Table 5: SpatialRGPT-Bench results. Accuracy (%) of Robot-R1, SFT baselines in evaluating the performance of VQA tasks from SpatialRGPT-Bench [20], which measures (a) how LVLMs accurately predict the positional relations between objects, and (b) how LVLMs precisely estimate the absolute object positions, where the prediction is considered correct if it lies within $\pm$25of the ground-truth values. In contrast to SFT models that struggle to generalize to novel spatial reasoning tasks, Robot-R1 effectively transfers to these spatial reasoning tasks.
(a) Quantitative results Model Direct Dist. Horiz. Dist. Vert. Dist. Width Height Direction Quant. Avg Qwen2.5-VL-7B-Ins [62] 1.35 9.02 8.49 6.77 9.77 28.04 9.88 + Direct SFT 4.05 3.28 0.94 0.75 4.51 42.06 8.41 + CoT SFT 1.35 2.46 0 3.01 6.77 38.32 7.88 + Robot-R1(Ours) 6.08 14.75 22.64 21.8 19.55 12.15 15.89 (b) Qualitative results Model Below/Above Left/Right Big/Small Tall/Short Wide/Thin Behind/Front Qual. Avg Qwen2.5-VL-7B-Ins [62] 35 50.48 6.6 17.86 27.88 36.36 29.07 + Direct SFT 3.33 17.14 0 9.82 0 0 5.02 + CoT SFT 15.83 33.33 45.28 39.28 47.12 35.45 35.62 + Robot-R1(Ours) 42.5 41.9 31.13 37.5 38.46 52.72 40.79
@@
Table 6: Ablation study. Evaluation results on the Robot-R1 Bench comparing different dataset designs, varying auxiliary tasks, i.e., waypoint prediction (WP), current state prediction (SP), and primitive movement prediction (MP), and the question format.
@@
QA Type Task Type Low Level Control High Level Control Open-End MCQA WP SP MP Movement Spatial Avg Planning Subtask Avg - 0.68 1.37 1.03 1.60 1.26 1.43 - 0.48 1.03 0.76 1.48 1.16 1.17 - 0.40 1.42 0.91 1.44 1.22 1.33 - 0.54 1.51 1.03 1.50 1.24 1.37 - 0.76 1.51 1.14 1.44 1.30 1.37
[…]
@@
Evolution of reasoning patterns during ROBOT-R1 training. We find that the evolution of reasoning patterns during Robot-R1 training diverges notably from trends reported in Deepseek-R1 [17]. While reasoning models for mathematical and coding tasks tend to develop longer reasoning chains over time, Robot-R1 exhibits the opposite pattern of producing progressively shorter and more focused reasoning during validation accuracy increase (see Figure 4). Qualitatively, the model initially generates broad action plans or subtask reasoning in an overview format. Over time, these are gradually replaced by more concise, logically structured reasoning traces that directly link relevant embodied reasoning components (see Appendix C for detailed examples).
((a)) Reward during Robot-R1 training
((b)) Response length during Robot-R1 training
Figure 4: Training metrics for Robot-R1 (a) Shows validation accuracy change over training steps. (b) Shows mean response length over training steps.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) have recently shown great promise in advancing robotics by combining embodied reasoning with robot control. A common approach involves training on embodied reasoning tasks related to robot control using Supervised Fine-Tuning (SFT). However, SFT datasets are often heuristically constructed and not explicitly optimized for improving robot control. Furthermore, SFT often leads to issues such as catastrophic forgetting and reduced generalization performance. To address these limitations, we introduce Robot-R1, a novel framework that leverages reinforcement learning to enhance embodied reasoning specifically for robot control. Robot-R1 learns to predict the next keypoint state required for task completion, conditioned on the current scene image and environment metadata derived from expert demonstrations. Inspired by the DeepSeek-R1 learning approach, Robot-R1 samples reasoning-based responses and reinforces those that lead to more accurate predictions. To rigorously evaluate Robot-R1, we also introduce a new benchmark that demands the diverse embodied reasoning capabilities for the task. Our experiments show that models trained with Robot-R1 outperform SFT methods on embodied reasoning tasks. Despite having only 7B parameters, Robot-R1 even surpasses GPT-4o on reasoning tasks related to low-level action control, such as spatial and movement reasoning.
**footnotetext: Equal advising
@@ 1 Introduction
Figure 1: Illustration of the Robot-R1 framework. (a) Robot-R1 uses robot states and image observations from expert demonstrations to create a dataset. (b) These data are reformulated into three different multiple-choice question answering (MCQA) tasks: predicting next states, current states, and movements. (c) During training, an LVLM solves MCQA tasks with reasoning which is then optimized using the GRPO algorithm [17] to reinforce reasoning pathways.
@@
The key idea of Robot-R1 is to train LVLMs to predict the next keypoint state necessary for task completion through a reasoning process, and to optimize this reasoning via RL to maximize prediction accuracy. However, since the keypoint lies in a continuous space, the action space is vast, making exploration particularly challenging. This makes it difficult for the model to efficiently learn reasoning strategies through trial and error. To address this, Robot-R1 reformulates the next-state prediction problem as a multiple-choice question-answering (QA) problem. This discrete formulation narrows the action space, making the learning process more efficient. To further enhance state understanding, we add two auxiliary QA tasks: (i) current state prediction QA, where the model predicts the robot’s current state from visual observations, and (ii) movement prediction QA, where the model predicts rule-based linguistic descriptions of state changes (e.g., “move up”, “move down”).
To analyze how Robot-R1 improves embodied reasoning capabilities in trained models, we design a novel benchmark called the Robot-R1 Bench (see Figure 2). Through this benchmark, we observe that models trained with Robot-R1 achieve over a 28% improvement in embodied reasoning tailored for low-level action control. In particular, despite having only the 7B parameters, Robot-R1 outperforms several major commercial models, including GPT-4o [25], in this domain. To evaluate whether the embodied reasoning abilities acquired through Robot-R1 transfer effectively to other tasks, we conduct evaluations on two external benchmarks. First, on EmbodiedBench Manipulation [59], a vision-driven robot agent benchmark, Robot-R1 yields a 31% improvement in task performance. Moreover, on SpatialRGPTbench [12], which tests 3D spatial reasoning, Robot-R1 achieves approximately 40% improvement in quantitative metrics and about 60% improvement in qualitative metrics. These results contrast with conventional SFT approaches, which show limited improvement in embodied reasoning and tend to exhibit performance degradation when applied to out-of-distribution tasks. This suggests that Robot-R1 can effectively learn generalizable and diverse embodied reasoning capabilities simply by next-state prediction.
@@ 3.1 Preliminaries
Group relative policy optimization. We employ the Group Relative Policy Optimization (GRPO) algorithm [17] to optimize the policy. Specifically, for a given query $q$, we generate $G$ different responses $\mathbf{o}=\{o_{1},o_{2},...,o_{G}\}$ from the current policy $\pi_{\theta_{\text{old}}}$. Each response $o_{i}$ is evaluated using the reward model to produce a corresponding reward $r_{i}$, resulting in the set $\mathbf{r}=\{r_{1},r_{2},...,r_{G}\}$. For each response, we compute an advantage score $A_{i}$, using the mean and standard deviation of the rewards for that query, i.e., $A_{i}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$. Finally, GRPO updates the policy by maximizing this advantage while applying a KL penalty, through the following objective:
@@ 3.2 Data Generation
Movement prediction QA. The goal of this task is to predict the movement (e.g., “move up”, “move right”) that the robot should take to move from the current state toward the next waypoint. To extract labels for these movement, we apply rule-based heuristics based on the change in 3D Cartesian space. Specifically, we compute the difference in position between the current and next waypoint states, $s_{k^{*}}-s_{t}$, and identify the primary direction of movement along the $x$, $y$, and $z$ axes. If there is a non-zero change along an axis, a corresponding directional command is generated (e.g., “move forward“ for a positive change in the x-axis); this label is denoted as $a_{t}$. If the change along an axis is less than or equal to half of the largest change across all axes, the adverb “slightly” is added (e.g., “Slightly move backward”). To construct a multiple-choice question, we randomly sample three adjacent movement $\{a^{\prime}_{d1},a^{\prime}_{d2},a^{\prime}_{d3}\}$. The question is formulated as $Q_{\tt{movement}}(M,o_{t},s_{t},\texttt{shuffle}(\{a_{t},a^{\prime}_{d1},a^{\prime}_{d2},a^{\prime}_{d3}\}))$ where the answer is $A=a_{t}$.
@@ 3.4 Robot-R1 Bench: A Novel Benchmark for Evaluating Embodied Reasoning
Existing visual question answering (VQA) benchmarks for embodied reasoning primarily focus on general visual understanding, without explicitly evaluating the nuanced reasoning processes behind robotic behavior [42, 57, 14]. Meanwhile, robotics-focused benchmarks often rely on simulation-based evaluation [59, 68, 36], prioritize task successes as the main metric, or focus on tasks that are not aligned with the table-top manipulation scenarios central to our work. While these benchmarks provide valuable insights, they do not directly assess a model’s ability to reason about robot actions. To address this gap, we introduce Robot-R1 Bench, a new benchmark designed to evaluate embodied reasoning through open-ended answering grounded in robot demonstrations.
Design features of Robot-R1 Bench. In practical applications, models typically face open-ended decisions rather than selecting from pre-defined options. To better reflect this, Robot-R1 Bench adopts an open-ended QA format that more closely resembles practical use cases. All questions are based on images from expert demonstrations, ensuring visually grounded and realistic robotic scenarios. Robot-R1 Bench supports fine-grained evaluation across four key reasoning types: planning, high-level action reasoning, movement reasoning, and spatial reasoning. Each question is crafted to assess one of these abilities, allowing for evaluation across both high-level decision making (planning and high-level action reasoning) and low-level control reasoning (movement reasoning and spatial reasoning).
Dataset construction. The Robot-R1 Bench dataset consists of 10 tasks from RLBench [27]. For each task, we randomly sample five frames from expert demonstrations, resulting in a total of 50 images. For each image, experienced researchers manually created questions and detailed reference answers, targeting the four key reasoning abilities. The final dataset consists of 215 open-ended questions in total: 65 for spatial reasoning and 50 for each of the other three reasoning types.
@@ 4 Experiment
4.1 Experimental Setup
Table 2: Summary of Robot-R1 Bench results. In embodied reasoning tailored for high-level control, Robot-R1 significantly outperforms SFT, the baseline training method.
Model Planning High-level Action GPT-4o [25] 1.96 2.02 Cluade-3-7-sonnet [5] 1.72 1.58 Gemini-2.0-flash [18] 1.84 1.12 Qwen2.5-VL-7B-Ins [7] 1.66 1.04 w/ Direct SFT 0 0.04 w/ CoT SFT 0.60 1.28 w/ Robot-R1 (Ours) 1.44 1.30
Training data. The training data used in our experiments are generated using the built-in data generator in RLBench [27]. We collect 50 demonstrations per task from the variation 0 settings, rendering them at 224 × 224 resolution using OpenGL3. Waypoints from these demonstrations are extracted using the waypoint extraction functionality provided in the ARM [26] repository [^1] https://github.com/stepjam/ARM. The tasks selected from RLBench for training our model include: pick_up_cup, push_button, put_rubbish_in_bin, phone_on_base, and take_lid_off_saucepan. For generating the waypoint prediction QA training data, as described in Section 3, a frame extraction interval of $t=10$ was used between the current frame ($o_{t},s_{t}$) and the selected future keypoint frame ($s_{k^{*}}$), similar to demonstration augmentation proposed in ARM. Consequently, each task contains approximately 2.5K questions, resulting in a total of around 7.5K QA pairs across the three QA tasks used for training.
@@ 4.1 Experimental Setup
Baseline models. The proposed Robot-R1 learns to predict the next waypoint state from expert demonstrations by generating an explicit reasoning process. To evaluate the importance of this learned reasoning, we establish two baseline models trained via Supervised Fine-tuning (SFT):
@@
    CoT SFT: This model is fine-tuned on a QA dataset where the input prompt includes a manually structured Chain-of-Thought (CoT) path. This CoT path sequentially incorporates planning, high-level action reasoning, and movement components to guide the prediction of the next waypoint state, similar to Zawalski et al. [64]. This baseline enables comparison between the reasoning learned via Robot-R1 against reasoning guided by a predefined, structured thought process.
@@
Implementation detail. Robot-R1 is trained using GRPO as described in Section 3. We use the hyperparameter configurations in the EasyR1 workspace. [^2] https://github.com/hiyouga/EasyR1 For the training process, we utilize a batch size of 128 over 5 epoch. During GRPO updates, 5 samples were generated per prompt with a sampling temperature of 1.0. The rollout batch size is set to 512. We use a learning rate of $1.0\times 10^{-6}$ with a weight decay of $1.0\times 10^{-2}$. For the SFT baselines, we use the same batch size, but learning rate is $1.0\times 10^{-5}$. We use the hyperparameters provided in the Qwen2-VL-Finetune workspace. [^3] https://github.com/2U1/Qwen2-VL-Finetune
@@ 4.2 Robot-R1 Bench
Results. Table 3 presents the results across the reasoning benchmarks focused on low-level control, and Figure 3 provides its visualization to help understanding. We find that commercial model performance improves consistently with their version updates, indicating the benchmark’s sensitivity to model quality. Notably, the Robot-R1-trained model achieves the highest performance across both the overall benchmark and its individual components, outperforming all commercial models evaluated. Specifically, in movement prediction, our method outperforms even the SFT-trained model, which is explicitly trained to predict accurate actions. Table 2 further shows the performance in reasoning tasks tailored for high-level control. Here, the correlation between model version and benchmark performance is less pronounced. We hypothesize that this is due to the inherently high variability in valid high-level actions for robotic control. Nevertheless, we still observe a general trend of increasing benchmark performance as model quality improves. Importantly, our Robot-R1 exhibits significant performance gains in high-level action reasoning compared to the base and the SFT-trained model. This suggests that effective high-level reasoning can emerge from training exclusively on low-level control information, even without explicit high-level action supervision. Interestingly, we find that performance on planning slightly decreases, likely because our training objective focuses primarily on next keypoint prediction rather than long-horizon planning.
@@ 4.3 EmbodiedBench Manipulation
w Direct SFT 0 0 0 0 0 0
w CoT SFT 0 0 0 0 0 0
w Robot-R1 (Ours) 12.5 8.3 6.3 14.6 16.7 11.68
@@
Setup. We evaluate the model performance on four different tasks: Pick & Place Objects, Stack Objects, Shape Sorter Placement, and Table Wiping. To comprehensively assess instruction-following capabilities, the benchmark categorizes task instructions into five groups: (i) Base, which requires fundamental task-solving skills, (ii) Common Sense, which requires general world knowledge to resolve indirect objective references, (iii) Complex instructions, which involves longer, possibly distracting context that obscures the primary command, (iv) Spatial reasoning, where objects are referred to by their spatial relationships to other objects, and (v) Visual understanding, which requires recognizing objects based on visual attributes such as color or shape. All models are evaluated with a sampling temperature of 0. For comparison, the experimental results of GPT-4o-mini [25] and Gemini-1.5-flash [48] are adopted directly from the values reported in their paper.
@@ 4.4 Ablation Study and Analysis
Table 5: SpatialRGPT-Bench results. Accuracy (%) of Robot-R1, SFT baselines in evaluating the performance of VQA tasks from SpatialRGPT-Bench [12], which measures (a) how LVLMs accurately predict the positional relations between objects, and (b) how LVLMs precisely estimate the absolute object positions, where the prediction is considered correct if it lies within $\pm$25% of the ground-truth values. In contrast to SFT models that struggle to generalize to novel spatial reasoning tasks, Robot-R1 effectively transfers to these spatial reasoning tasks.
(a) Quantitative results Model Direct Dist. Horiz. Dist. Vert. Dist. Width Height Direction Quant. Avg Qwen2.5-VL-7B-Ins [7] 1.35 9.02 8.49 6.77 9.77 28.04 9.88 w Direct SFT 4.05 3.28 0.94 0.75 4.51 42.06 8.41 w CoT SFT 1.35 2.46 0 3.01 6.77 38.32 7.88 w Robot-R1 (Ours) 6.08 14.75 22.64 21.8 19.55 12.15 15.89 (b) Qualitative results Model Below/Above Left/Right Big/Small Tall/Short Wide/Thin Behind/Front Qual. Avg Qwen2.5-VL-7B-Ins [7] 35 50.48 6.6 17.86 27.88 36.36 29.07 w Direct SFT 3.33 17.14 0 9.82 0 0 5.02 w CoT SFT 15.83 33.33 45.28 39.28 47.12 35.45 35.62 w Robot-R1 (Ours) 42.5 41.9 31.13 37.5 38.46 52.72 40.79
@@
Table 6: Pearson correlation between human and GPT-4o [25] judgments on Robot-R1 Bench. Computed over 80 responses (from Robot-R1 and GPT-4o), where human scores are defined as the median of 8 human expert annotations.
Planning High-level Action Movement Spatial
Pearson Correlation 0.3315 0.8974 0.8931 0.8961
Validation of the Robot-R1 bench. To confirm the validity of the Robot-R1 Bench’s LLM-as-judge, we measure the Pearson correlation between LLM and human assigned score. Considering the cost for collecting human annotations, we sample 40 problems from the Robot-R1 Bench and collect judgments from 8 participants with expertise in robotics. The experiment is conducted as a blind evaluation of responses produced by Robot-R1 and GPT-4o on a total of 80 problems. We define the human score as the median across the 8 participants. As shown in Table 6, the Pearson correlations between LLM and human scores are close to 0.9 for all tasks except planning, supporting the reliability of Robot-R1 Bench. For the planning task, we observe a lower correlation, likely because its longer and more open-ended responses allow multiple valid solutions, making precise automatic judging inherently more challenging. Notably, the tasks with higher correlations, such as high-level action, movement and spatial, are also where Robot-R1 achieves the largest performance gains, further supporting the validity of the improvements achieved by Robot-R1.
Table 7: Robot-R1 Bench results on different RL algorithms.
Model Planning High-level Action Movement Spatial GRPO [43] 1.44 1.30 0.76 1.51 RLOO [1] 1.56 1.54 0.68 1.52 REINFORCE++ [19] 1.40 1.08 0.62 1.38
Effect of different RL algorithms. We evaluate the effectiveness of GRPO used to train Robot-R1 by training additional models with RLOO [1] and REINFORCE++ [19], two RL algorithms commonly employed for LLM optimization. We use the default hyperparameters provided by the workspace  [^4] https://github.com/hiyouga/EasyR1 and evaluate all models on the Robot-R1 Bench. Table 7 shows that RLOO achieves performance comparable to GRPO; however, REINFORCE++ performs worse on average across tasks. We hypothesize that this is due to weaker reward normalization. Specifically, REINFORCE++ performs batch-level reward normalization, which often induces higher variance than the prompt-query-level normalization applied in GRPO and RLOO, leading to less stable learning and limited performance improvements. These findings indicate that RL algorithms that incorporate variance-reducing mechanisms, such as GRPO or RLOO, are beneficial for Robot-R1.
Table 8: Robot-R1 Bench results across random seeds. Robot-R1 consistently improves performance over the non-finetuned original model (Qwen2.5-VL-7B-Ins [7]) across different random seeds, showing strong robustness to seed variability.
Model Planning High-level Action Movement Spatial
Qwen2.5-VL-7B-Ins [7] 1.66 1.04 0.58 1.40
w Robot-R1 (Ours) 1st Seed 1.44 1.30 0.76 1.51
w Robot-R1 (Ours) 2nd Seed 1.70 1.46 0.64 1.66
w Robot-R1 (Ours) 3rd Seed 1.56 1.50 0.54 1.60
Seed Avg. 1.57 1.42 0.65 1.59
Seed Std. 0.13 0.11 0.11 0.08
Robustness across random seeds. To evaluate whether Robot-R1 consistently yields performance gains across different random seeds, we conduct two additional experimental runs on the same dataset while changing only the random seed for online-sampling. We then evaluate all models on the Robot-R1 Bench. As shown in Table 8, Robot-R1 consistently achieves higher performance than non-finetuned model across all seeds. The improvements remain significant when accounting for variance. These results demonstrate that training with Robot-R1 is robust to seed variability.
@@
Table 9: Ablation study. Evaluation results on the Robot-R1 Bench comparing different dataset designs, varying auxiliary tasks, i.e., waypoint prediction (WP), current state prediction (SP), and movement prediction (MP), and the question format.
@@
QA Type Task Type Low Level Control High Level Control Open-End MCQA WP SP MP Movement Spatial Avg Planning High-level Action Avg - 0.68 1.37 1.03 1.60 1.26 1.43 - 0.48 1.03 0.76 1.48 1.16 1.17 - 0.40 1.42 0.91 1.44 1.22 1.33 - 0.54 1.51 1.03 1.50 1.24 1.37 - 0.76 1.51 1.14 1.44 1.30 1.37
Impact of QA design on performance. Table 9 demonstrates the effects of different QA design choices on model performance. Our results show that adding auxiliary tasks progressively improves overall performance. We also evaluate an alternative approach that uses open-ended generation instead of multiple-choice question answering (MCQA). In this setup, the model directly generates next waypoint values in an open-ended format and uses $\text{clip}(1-\text{L1},0,1)$ as the reward function, where the L1 represents the distance between the predicted and ground-truth states. While the open-ended approach initially outperforms single-task MCQA, it exhibits significant performance degradation when auxiliary tasks – such as current-state prediction – are introduced. Consequently, the MCQA framework demonstrates superior results. These findings highlight the critical role of question structure design in reinforcement learning for embodied reasoning.
@@
Qualitative analysis. We find that the evolution of reasoning patterns during Robot-R1 training diverges notably from trends reported in Deepseek-R1 [17]. While reasoning models for mathematical and coding tasks tend to develop longer reasoning chains over time, Robot-R1 exhibits the opposite pattern of producing progressively shorter and more focused reasoning (see Appendix D for detailed examples). Qualitatively, the model initially generates broad planning or high-level action reasoning in an overview format. Over time, these are gradually replaced by more concise, logically structured reasoning traces that directly link relevant embodied reasoning components.